# Dongseok Kim    Hyoungsun Choi    Mohamed Jismy Aashik Rasool    Gisung Oh

### v1

@@ 2 Related Work
Several works lift instance-wise attributions to global additive descriptions of a predictor by decomposing $f$ into low-dimensional main and interaction components and showing that tools such as partial dependence and interventional Shapley values arise as linear functionals of this decomposition (hiabu2023unifying; fumagalli2024unifying). Other lines aggregate variable-importance measures or local explanations across Rashomon sets, datasets, or examples, treating them as distributions or geometric objects in an ”explanation space” and using them to explain distribution shifts or recourse patterns rather than a single model-specific ranking (donnelly2023rashomon; babbar2024different; xenopoulos2022gale; alfeo2023local). These methods demonstrate that local explanations can be turned into global effect summaries, but they typically provide point estimates of importance and do not attach selection-adjusted uncertainty to feature-wise statistics.
@@
A line of work formalizes global variable importance as a population-level contrast in oracle predictiveness between models that use all features and models that exclude a given subset, and derives efficient estimators whose asymptotic normality enables confidence intervals and hypothesis tests even with flexible learners (williamson2023general; williamson2020efficient). Building on this perspective, decorrelated importance parameters mitigate the impact of feature correlation while still supporting valid uncertainty quantification in high-dimensional or genomics applications via Shapley-style loss decompositions and bootstrap procedures (verdinelli2024decorrelated; johnsen2023inferring). Additional methods develop minipatch ensembles, simultaneous confidence sets for feature rankings, and conditional permutation schemes with formal type I error control for variable inclusion decisions (gan2022model; neuhof2024confident; chamma2023statistically). These approaches provide principled uncertainty quantification for pre-specified global importance functionals, but they do not explicitly model a feature-selection step or adjust their intervals and $p$-values for data-driven selection events.
@@
Shapley-based attributions have been linked to conditional independence testing, for example in SHAP-XRT, which embeds local Shapley scores into a conditional randomization framework to obtain finite-sample valid tests for feature relevance (teneggi2022shap). Other approaches design attribution-driven screening rules in which Shapley scores are compared against data-driven thresholds or causal criteria, yielding adaptive selection procedures such as ShapHT plus and Causal SHAP that control false discoveries or preserve causal interpretability (yin2022adaptive; lamsaf2025causal). Complementary work studies the statistical significance of entire attribution rankings, proposing multiple-hypothesis tests that retrospectively certify the correctness of the top-$K$ features identified by an importance estimator (goldwasser2024statistical). These lines of work use attribution scores as inputs to hypothesis tests or screening rules, but they typically focus on local conditional independence decisions or on validating a fixed ranking, rather than producing a global feature-importance table with post-selection uncertainty for each selected feature.
@@
Several works modify local post hoc explainers to improve the stability of attributions under small input perturbations, for example by redesigning the sampling and weighting scheme of LIME in GLIME or by leveraging invariant risk minimization ideas to obtain explanations that vary smoothly across nearby examples (tan2023glime; dhurandhar2023locally). Other methods explicitly target uncertainty and stability in explanations, introducing Bayesian variants of attribution methods with credible intervals, learning global surrogate explanations that remain faithful under shifts, and designing rule-based or counterfactual objectives and evaluation metrics for stability in tabular decision problems (slack2021reliable; lakkaraju2020robust; guidotti2024stable; sepulveda2025enhancing). Cross-model robustness is studied by aggregating explanations across committees of models and using cross-model consensus scores to capture features that are relied on consistently across architectures (li2023cross). These efforts highlight the importance of stability and cross-model agreement, which we also target in our empirical evaluation, but they do not provide selective-inference-style uncertainty for feature-wise effects.
@@
The post-selection inference literature develops tools for valid regression inference after data-driven feature selection by conditioning on the selection event, often exploiting the fact that procedures such as the lasso, forward stepwise, and least-angle regression induce polyhedral constraints on the response so that linear contrasts of the coefficients follow truncated normal laws (lee2016exact; tibshirani2016exact). Within a more general exponential-family framework, selective tests and confidence intervals are constructed to be most powerful and unbiased under arbitrary model-selection rules, providing a principled alternative to naive plug-in $t$-tests that ignore selection (fithian2014optimal; taylor2015statistical). Subsequent extensions introduce randomized selection steps, approximate Bayesian selective-likelihood formulations, and knockoff filters that address high-dimensional variable selection via synthetic controls and finite-sample false-discovery guarantees (tian2018selective; panigrahi2021integrative; barber2019knockoff). Together, these approaches show how to attach valid $p$-values and confidence intervals to coefficients in regression models chosen by adaptive feature selection, but they are typically developed for generic supervised learning tasks and do not directly address Shapley-guided global feature-importance tables for black-box predictors.
@@ 3 Method
Given $(f,D)$, the $\phi$-test produces a global feature-importance table whose rows index features $j=1,\dots,p$ and whose columns report: (i) a SHAP-based global score $I_{j}$; (ii) a surrogate regression coefficient $\widehat{\beta}_{j}$; and (iii) post-selection $p$-values and confidence intervals for features retained after a SHAP-guided selection step. Algorithm 1 summarizes the procedure.
@@ 3.3 Stage 3: Selective inference for surrogate coefficients
whose conditional law under the polyhedral selection event $\mathcal{E}$ is univariate Gaussian truncated to an interval determined by $(A,b)$. Using this truncated normal law with $\sigma^{2}$ replaced by $\widehat{\sigma}^{2}$, we compute a selective $p$-value and confidence interval for $\beta_{j}$ and write
@@ 3.4 Global feature-importance table
The $\phi$-test outputs a table indexed by features $j=1,\dots,p$. For each feature we report: (i) the SHAP-based global score $I_{j}$; (ii) the surrogate coefficient estimate $\widehat{\beta}_{j}$ if $j\in S$ (blank otherwise); and (iii) the selective $p$-value $p_{j}^{\mathrm{sel}}$ and confidence interval $\mathrm{CI}_{j}$ if $j\in S$. Features with large $I_{j}$ but $j\notin S$ are globally influential under SHAP yet not supported by the surrogate at the chosen selection threshold, whereas features in $S$ with modest $I_{j}$ correspond to more subtle but statistically stable effects. By presenting these quantities in a single table, the $\phi$-test provides an ablation-free, global explanation of how the black-box depends on its inputs and attaches post-selection uncertainty to each selected feature.
@@
9:  $(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j})\leftarrow\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr)$
@@ 4 Theory
This section formalizes the statistical guarantees underlying the $\phi$-test. We work under a Gaussian surrogate model for the black-box predictions and a polyhedral representation of the selection event, and use standard selective-inference tools to characterize the conditional law of linear contrasts. Throughout, the black-box predictor $f$ and covariates $X$ are treated as fixed, and randomness arises only from the surrogate response.
@@ 4.1 Selective-inference setup
denote the ordinary least-squares estimator in the surrogate model restricted to $S$. For each $j\in S$, we are interested in testing
$H_{0,j}:\beta_{j}=0$
and constructing a confidence interval for $\beta_{j}$, conditional on the event $\mathcal{E}$ that $S$ has been selected.
@@ 4.2 Truncated-normal law and selective validity
is proportional to $\widehat{\beta}_{j}$; for example, when $S$ is fixed one may take
$v_{j}=e_{j}^{\top}(X_{S}^{\top}X_{S})^{-1}X_{S}^{\top},$
where $e_{j}$ is the canonical basis vector for feature $j$ within $S$. Writing $\mu=X_{S_{0}}\beta$ in (1), define
@@
Let $S_{0}=\mathcal{T}_{M}\bigl((I_{j})_{j=1}^{p}\bigr)$ denote the candidate feature set produced by Stage 1 (Subsection 3.1) of the $\phi$-test based on global SHAP scores. Under (1), the set $S_{0}$ is non-random with respect to $y^{(\mathrm{sur})}$, so all probabilities can be interpreted conditionally on $(f,X,S_{0})$.
@@
Work under (1) and (2), and condition on $(f,X,S_{0})$ as in Lemma 4.1. Fix $j\in S$ and consider $H_{0,j}:\beta_{j}=0$. Let $p_{j}^{\mathrm{sel}}$ and $\mathrm{CI}_{j}$ denote the selective $p$-value and $(1-\alpha)$ confidence interval returned by $\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr)$. Then, for all $\alpha\in[0,1]$,
@@
$\mathbb{P}\bigl(\beta_{j}\in\mathrm{CI}_{j}\,\big|\,f,X,S_{0},j\in S\bigr)$ $\geq 1-\alpha.$
That is, the selective $p$-values control conditional type-I error and the intervals achieve at least nominal conditional coverage for the surrogate coefficients given that feature $j$ has been selected.
@@ 5.2 Compared methods
  • SHAP-TopK. (lundberg2017unified) Rank features by $I_{j}=\frac{1}{n}\sum_{i=1}^{n}|\phi_{j}(x_{i})|$ and select the top $K$ features.
  • SPVIM-Boot. (williamson2020efficient) Shapley-based variable importance with bootstrap confidence intervals; select features whose intervals exclude zero.
  • SHAP-HT. (yin2022adaptive) Marginal SHAP-based hypothesis tests of feature relevance with multiple-testing correction.
  • StableSHAP. (goldwasser2024statistical) A stability-oriented aggregation that re-fits the backbone on bootstrap samples and selects features that appear frequently among the top-ranked SHAP features.
@@ 5.3 Evaluation criteria
We report four scalar criteria for performance, sparsity, stability, and cross-backbone robustness. Let $\hat{S}$ denote the selected feature set and let $\mathrm{Perf}(\cdot)$ be a test performance metric (we use $R^{2}$ in all experiments) of a linear model fitted using only features in $\hat{S}$.
@@ 5.4 Quantitative results
5.5 Example global selective inference table
Table 5 shows the global feature-importance table produced by $\phi$-test on AirQuality with an XGBoost backbone. For the strongest pollutants, SHAP scores, regression coefficients, and very small selective $p$-values all indicate a positive and statistically well-supported effect, with selective 95% confidence intervals lying strictly above zero. In contrast, NO2(GT) has a moderate SHAP score but a selective $p$-value of $0.12$ and a confidence interval that crosses zero, suggesting that its effect is not conclusively different from zero once conditioning on the selection event. This illustrates how the table combines Shapley-based global importance with post-selection $p$-values and intervals to distinguish strongly supported features from borderline ones.
Table 5: Split-sample selective regression table for AirQuality ($\phi$-test, XGBoost backbone). Reported $p$-values and confidence intervals come from $t$-tests on the held-out inference split.
Feature SHAP Coef SE $z$ p-value 95% CI
@@ Appendix A Notation for Algorithm 1
$\mathcal{I}_{\mathrm{TN}}(T_{j};A,b,\widehat{\sigma}^{2})$ Truncated-normal map returning $(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j})$
@@ B.1 Truncated-normal law for linear contrasts
used in the main text: under $H_{0,j}:\theta_{j}=0$ we evaluate tail probabilities in the truncated-normal law of $T_{j}$ conditional on $(W_{j},\mathcal{E})$, and we invert this family of tests to obtain selective confidence intervals for $\beta_{j}$.
@@ B.2 Proof of selective validity
$\mathbb{P}\bigl(\beta_{j}\in\mathrm{CI}_{j}\,\big|\,f,X,S_{0},j\in S\bigr)\geq 1-\alpha,$
where we use that $\theta_{j}=v_{j}^{\top}X_{S_{0}}\beta$ is the linear functional of $\beta$ corresponding to feature $j$ in the surrogate model. This establishes the conditional coverage claim in Proposition 4.2.

### v2

@@ 2 Related Work
Several works lift instance-wise attributions to global additive descriptions of a predictor by decomposing $f$ into low-dimensional main and interaction components and showing that tools such as partial dependence and interventional Shapley values arise as linear functionals of this decomposition (Hiabu et al., 2023; Fumagalli et al., 2024). Other lines aggregate variable-importance measures or local explanations across Rashomon sets, datasets, or examples, treating them as distributions or geometric objects in an ”explanation space” and using them to explain distribution shifts or recourse patterns rather than a single model-specific ranking (Donnelly et al., 2023; Babbar et al., 2024; Xenopoulos et al., 2022; Alfeo et al., 2023). These methods demonstrate that local explanations can be turned into global effect summaries, but they typically provide point estimates of importance and do not attach selection-adjusted uncertainty to feature-wise statistics.
@@
A line of work formalizes global variable importance as a population-level contrast in oracle predictiveness between models that use all features and models that exclude a given subset, and derives efficient estimators whose asymptotic normality enables confidence intervals and hypothesis tests even with flexible learners (Williamson et al., 2023; Williamson and Feng, 2020). Building on this perspective, decorrelated importance parameters mitigate the impact of feature correlation while still supporting valid uncertainty quantification in high-dimensional or genomics applications via Shapley-style loss decompositions and bootstrap procedures (Verdinelli and Wasserman, 2024; Johnsen et al., 2023). Additional methods develop minipatch ensembles, simultaneous confidence sets for feature rankings, and conditional permutation schemes with formal type I error control for variable inclusion decisions (Gan et al., 2022; Neuhof and Benjamini, 2024; Chamma et al., 2023). These approaches provide principled uncertainty quantification for pre-specified global importance functionals, but they do not explicitly model a feature-selection step or adjust their intervals and $p$-values for data-driven selection events.
@@
Shapley-based attributions have been linked to conditional independence testing, for example in SHAP-XRT, which embeds local Shapley scores into a conditional randomization framework to obtain finite-sample valid tests for feature relevance (Teneggi et al., 2022). Other approaches design attribution-driven screening rules in which Shapley scores are compared against data-driven thresholds or causal criteria, yielding adaptive selection procedures such as ShapHT plus and Causal SHAP that control false discoveries or preserve causal interpretability (Yin et al., 2022; Lamsaf et al., 2025). Complementary work studies the statistical significance of entire attribution rankings, proposing multiple-hypothesis tests that retrospectively certify the correctness of the top-$K$ features identified by an importance estimator (Goldwasser and Hooker, 2024). These lines of work use attribution scores as inputs to hypothesis tests or screening rules, but they typically focus on local conditional independence decisions or on validating a fixed ranking, rather than producing a global feature-importance table with post-selection uncertainty for each selected feature.
@@
Several works modify local post hoc explainers to improve the stability of attributions under small input perturbations, for example by redesigning the sampling and weighting scheme of LIME in GLIME or by leveraging invariant risk minimization ideas to obtain explanations that vary smoothly across nearby examples (Tan et al., 2023; Dhurandhar et al., 2023). Other methods explicitly target uncertainty and stability in explanations, introducing Bayesian variants of attribution methods with credible intervals, learning global surrogate explanations that remain faithful under shifts, and designing rule-based or counterfactual objectives and evaluation metrics for stability in tabular decision problems (Slack et al., 2021; Lakkaraju et al., 2020; Guidotti et al., 2024; Sepulveda et al., 2025). Cross-model robustness is studied by aggregating explanations across committees of models and using cross-model consensus scores to capture features that are relied on consistently across architectures (Li et al., 2023). These efforts highlight the importance of stability and cross-model agreement, which we also target in our empirical evaluation, but they do not provide selective-inference-style uncertainty for feature-wise effects.
@@
The post-selection inference literature develops tools for valid regression inference after data-driven feature selection by conditioning on the selection event, often exploiting the fact that procedures such as the lasso, forward stepwise, and least-angle regression induce polyhedral constraints on the response so that linear contrasts of the coefficients follow truncated normal laws (Lee et al., 2016; Tibshirani et al., 2016). Within a more general exponential-family framework, selective tests and confidence intervals are constructed to be most powerful and unbiased under arbitrary model-selection rules, providing a principled alternative to naive plug-in $t$-tests that ignore selection (Fithian et al., 2014; Taylor and Tibshirani, 2015). Subsequent extensions introduce randomized selection steps, approximate Bayesian selective-likelihood formulations, and knockoff filters that address high-dimensional variable selection via synthetic controls and finite-sample false-discovery guarantees (Tian and Taylor, 2018; Panigrahi et al., 2021; Barber and Candès, 2019). Together, these approaches show how to attach valid $p$-values and confidence intervals to coefficients in regression models chosen by adaptive feature selection, but they are typically developed for generic supervised learning tasks and do not directly address Shapley-guided global feature-importance tables for black-box predictors.
@@ 3 Method
Given $(f,\mathcal{D})$, the $\phi$-test produces a global feature-importance table whose rows index features $j=1,\dots,p$; for presentation, we may additionally append a residual (unselected) row that aggregates the SHAP mass of features not retained after selection. The table’s columns report: (i) a SHAP-based global score $I_{j}$; (ii) a surrogate regression coefficient $\widehat{\beta}_{j}$; and (iii) $p$-values and confidence intervals for retained features, computed either on a held-out split or via truncated-normal selective inference. Algorithm 1 summarizes the procedure.
@@ 3.3 Stage 3: Selective inference for surrogate coefficients
whose conditional law under the polyhedral selection event $\mathcal{E}$ is univariate Gaussian truncated to an interval determined by $(A,b)$. Using this truncated normal law with $\sigma^{2}$ replaced by $\widehat{\sigma}^{2}$, we compute a selective $p$-value and confidence interval for $\theta_{j}$ (the linear-contrast surrogate target) and write
@@ 3.4 Global feature-importance table
The $\phi$-test outputs a table whose rows index features $j=1,\ldots,p$; optionally, we append a residual (unselected) row that aggregates the SHAP mass of features not retained after selection. For each feature we report: (i) the SHAP-based global score $I_{j}$; (ii) the surrogate coefficient estimate $\widehat{\beta}_{j}$ if $j\in S$ (blank otherwise); and (iii) the selective $p$-value $p_{j}^{\mathrm{sel}}$ and confidence interval $\mathrm{CI}_{j}$ if $j\in S$. Features with large $I_{j}$ but $j\notin S$ are globally influential under SHAP yet not supported by the surrogate at the chosen selection threshold, whereas features in $S$ with modest $I_{j}$ correspond to more subtle but statistically stable effects. By presenting these quantities in a single table, the $\phi$-test provides an ablation-free, global explanation of how the black-box depends on its inputs and attaches post-selection uncertainty to each selected feature.
@@
9:  $(p^{\mathrm{sel}}_{j},\mathrm{CI}_{j})\leftarrow\mathcal{I}\!\left(T_{j};A,b,\widehat{\sigma}^{2}\right)$
@@ 4 Theory
This section formalizes the statistical guarantees underlying the $\phi$-test. We work under a Gaussian surrogate model for the black-box predictions and a polyhedral representation of the selection event, and use standard selective-inference tools to characterize the conditional law of linear contrasts. Throughout, we treat $f$ and $X$ as fixed and adopt a working Gaussian model for $y^{(\mathrm{sur})}=f(X)$, with noise absorbing surrogate mismatch.
@@ 4.1 Selective-inference setup
denote the ordinary least-squares estimator in the surrogate model restricted to $S$. For each $j\in S$, define the contrast statistic
$T_{j}\;=\;v_{j}^{\top}y^{(\mathrm{sur})},\qquad\theta_{j}\;=\;v_{j}^{\top}\mu,\quad\text{where }\mu=X_{S_{0}}\beta,$
so that $T_{j}\sim\mathcal{N}(\theta_{j},\,\sigma^{2}\|v_{j}\|_{2}^{2})$ under (1). We are interested in testing
$H_{0,j}:\theta_{j}=0$
and constructing a confidence interval for $\theta_{j}$, conditional on the event $\mathcal{E}$ that $S$ has been selected.
@@ 4.2 Truncated-normal law and selective validity
is proportional to the refit OLS coefficient for feature $j$ after Stage 2. For example, when $S$ is fixed and $\widehat{\beta}_{S}$ denotes the OLS fit on $X_{S}$, one may take
$v_{j}^{\top}=e_{j}^{\top}(X_{S}^{\top}X_{S})^{-1}X_{S}^{\top},$
where $e_{j}$ is the canonical basis vector for feature $j$ within $S$. Writing $\mu=X_{S_{0}}\beta$ in (1), define the target
@@
Let $S_{0}=\mathcal{T}_{M}\bigl((I_{j})_{j=1}^{p}\bigr)$ denote the candidate feature set produced by Stage 1 (Subsection 3.1) of the $\phi$-test based on global SHAP scores. Under (1), and conditional on $(f,X)$, the set $S_{0}$ is measurable with respect to $(f,X)$ and hence does not depend on $y^{(\mathrm{sur})}$; therefore all probabilities can be interpreted conditionally on $(f,X,S_{0})$.
@@
Work under (1) and (2), and condition on $(f,X,S_{0})$ as in Lemma 4.1. Fix $j\in S$ and consider $H_{0,j}:\theta_{j}=0$. Let $p_{j}^{\mathrm{sel}}$ and $\mathrm{CI}_{j}$ denote the selective $p$-value and $(1-\alpha)$ confidence interval returned by $\mathcal{I}_{\mathrm{TN}}\!\bigl(T_{j};A,b,\widehat{\sigma}^{2}\bigr)$. Then, for all $\alpha\in[0,1]$,
@@
$\mathbb{P}\bigl(\theta_{j}\in\mathrm{CI}_{j}\,\big|\,f,X,S_{0},j\in S\bigr)$ $\geq 1-\alpha.$
That is, the selective $p$-values control conditional type-I error and the intervals achieve at least nominal conditional coverage for the surrogate target $\theta_{j}$ given that feature $j$ has been selected.
@@ 5.2 Compared methods
  • SHAP-TopK. (Lundberg and Lee, 2017) Rank features by $I_{j}=\frac{1}{n}\sum_{i=1}^{n}|\phi_{j}(x_{i})|$ and select the top $K$ features.
  • SPVIM-Boot. (Williamson and Feng, 2020) Shapley-based variable importance with bootstrap confidence intervals; select features whose intervals exclude zero.
  • SHAP-HT. (Yin et al., 2022) Marginal SHAP-based hypothesis tests of feature relevance with multiple-testing correction.
  • StableSHAP. (Goldwasser and Hooker, 2024) A stability-oriented aggregation that re-fits the backbone on bootstrap samples and selects features that appear frequently among the top-ranked SHAP features.
@@ 5.3 Evaluation criteria
We report four scalar criteria for performance, sparsity, stability, and cross-backbone robustness. Let $\hat{S}$ denote the selected feature set (for $\phi$-test, $\hat{S}:=S$ in Algorithm 1) and let $\mathrm{Perf}(\cdot)$ be a test performance metric (we use $R^{2}$ in all experiments) of a linear model fitted using only features in $\hat{S}$.
@@ 5.4 Quantitative results
5.5 Example global inference table
Table 5 shows the global feature-importance table produced by the $\phi$-test on AirQuality with an XGBoost backbone. Stage 1 reports SHAP-based global importance on the training split, while the reported coefficients, standard errors, $p$-values, and 95% confidence intervals are computed on a held-out inference split after refitting a linear surrogate. Because the inference split is independent of the feature-screening step, the reported $p$-values and intervals are valid without post-selection adjustment.
For the strongest pollutants, SHAP scores and the refit regression estimates both indicate positive effects, with 95% confidence intervals lying strictly above zero. In contrast, NO2(GT) has a moderate SHAP score but a $p$-value of $0.12$ and a confidence interval that crosses zero, suggesting that its effect is not conclusively different from zero on the inference split. This illustrates how the table combines Shapley-based global importance with out-of-sample regression inference to distinguish strongly supported features from borderline ones.
Table 5: Split-sample regression table for AirQuality ($\phi$-test, XGBoost backbone). Coefficients and standard errors come from an OLS refit on the held-out inference split; reported $t$-statistics, $p$-values, and 95% confidence intervals are from two-sided $t$-tests on that split.
Feature SHAP Coef SE $t$ p-value 95% CI
@@ Appendix A Notation for Algorithm 1
$\mathcal{I}_{\mathrm{TN}}(T_{j};A,b,\widehat{\sigma}^{2})$ Truncated-normal map returning $(p_{j}^{\mathrm{sel}},\mathrm{CI}_{j})$ (or classical $t$-test/CI under split-sample)
@@ B.1 Truncated-normal law for linear contrasts
used in the main text: under $H_{0,j}:\theta_{j}=0$ we evaluate tail probabilities in the truncated-normal law of $T_{j}$ conditional on $(W_{j},\mathcal{E})$, and we invert this family of tests to obtain selective confidence intervals for $\theta_{j}$.
@@ B.2 Proof of selective validity
$\mathbb{P}\bigl(\theta_{j}^{\star}\in\mathrm{CI}_{j}\,\big|\,f,X,S_{0},j\in S\bigr)\geq 1-\alpha,$
where $\theta_{j}^{\star}=v_{j}^{\top}\mu=v_{j}^{\top}X_{S_{0}}\beta$ is the linear contrast (surrogate target) corresponding to feature $j$. This establishes the conditional coverage claim in Proposition 4.2.